# Reconstructing and resampling: a guide to utilising posterior samples from gravitational wave observationsA

### v1

@@ 1 Introduction
Gravitational wave observations of merging black holes and neutron stars are now routinely made by interferometric gravitational wave detectors. There are currently four operational kilometre-scale interferometes: Laser Interferometer Gravitational-Wave Observatory (LIGO: Aasi et al., 2015), Virgo (Acernese et al., 2015), and KAGRA (Akutsu et al., 2021). Their raw output is a time series of calibrated strain, representing the relative difference in arm lengths between the instruments. The –Virgo–KAGRA (LVK) Collaborations publish this data (Abac et al., 2025a) in addition to the Gravitational-Wave Transient Catalog (GWTC: Abac et al., 2025c), containing a list of candidate signals and measurements of their Bayesian posterior distributions (Abac et al., 2025e).
The posterior distribution (see Thrane & Talbot, 2019, for a discussion on why a Bayesian framework is used) is defined as the probability density of a set of source parameters $\theta$ conditional on the data $d$ and model $M$ (which implicitly defines the parameters) and can be calculated up to a normalisation factor from
@@
$\mathcal{Z}(d|M)=\int\mathcal{L}(d;\theta,M)\pi(\theta|M)\,d\theta\,.$ (2)
@@
The LVK analysis methodology for approximating Eq.˜1 for observed signals is described in detail in Section 5 of Abac et al. (2025d). In brief, for transient gravitational-wave data analysis, analysis is performed on a time series of consecutive strain data observations which we denote by $d$ which has a duration $T$ and sampling frequency $f_{s}$. Inference is performed on the frequency series $\tilde{d}$ generated by applying a fast Fourier transform and dividing by $f_{s}$ (a normalisation choice; see Thrane & Talbot (2019)). For signals lasting up to a few hundred seconds, we assume the non-astrophysical noise to be generated by a stationary additive coloured Gaussian noise process with a power spectral density (PSD) $P$ such that the likelihood for the $j$th frequency bin is given by
$\mathcal{L}(d_{j}|\theta,d)=\frac{1}{2\pi P_{j}}\exp\left(-\frac{2}{T}\frac{|\tilde{d}_{j}-\tilde{\mu}_{j}(\theta)|^{2}}{P_{j}}\right)\,$, (3)
@@
where $\tilde{\mu}_{j}(\theta)$ is the frequency-domain astrophysical signal model projected onto the response of the detector. The full likelihood is then calculated from the sum in log-space of the log-likelihood over $d_{j}$.
@@
Typically, $\tilde{\mu}_{j}(\theta)$ is a waveform approximant that models the inspiral–merger–ringdown of the compact binary coalescence (CBC) signal utilising either phenomenological fits (Ajith et al., 2007), the effective one-body approach (Buonanno & Damour, 1999, 2000), or numerical-relativity surrogates (Blackman et al., 2017). To date, Bilby uses the LALSuite (LIGO Scientific Collaboration et al., 2018; Wette, 2020) library to access these waveforms.
@@
Meanwhile, the prior distribution $\pi(\theta|M)$ is either generally weakly astrophysically motivated or informed by the observed population of CBC sources (Abac et al., 2025d).
@@
Eq.˜1 is not solvable in closed form for the CBC likelihood and prior, and the distribution is typically non-Gaussian, multimodal, and contains strong curved degeneracies. Therefore, computational Bayesian inference methods such as Markov chain Monte Carlo (MCMC; Hastings, 1970) and nested sampling (Skilling, 2006) are used with post-processing to produce a set of equally-weighted posterior samples $\theta_{i}\sim p(\theta|d,M)$ and an estimate of the evidence $\mathcal{Z}(d|M)$. To date, the LVK has used LALInference (Veitch et al., 2015), RIFT (Pankow et al., 2015; Lange et al., 2017; Wysocki et al., 2019), and Bilby (Ashton et al., 2019; Romero-Shaw et al., 2020), with an adaptation of the Dynesty (Speagle, 2020) nested sampling implementation, to produce posterior samples. In this work, we will consider the LVK posteriors produced by Bilby and packaged using PESummary (Hoy & Raymond, 2021). This is motivated by the ease of use in reconstructing the posteriors using Bilby (part of its design philosophy) and the wide applicability: Bilby posteriors have been estimated for all signals above threshold for parameter estimation from the first observing run until the most recent release, see GWTC-2.1 (Abbott et al., 2024), GWTC-3.0 (Abbott et al., 2023), and GWTC-4.0 (Abac et al., 2025e).
@@
The posterior samples from a given observed signal form a data set that can be used for further analyses. For example, taking samples from a set of highly probable signals, the samples form an input data set that can be used to constrain the astrophysical distribution of merging black holes (see, e.g. Talbot & Thrane, 2018; Abac et al., 2025f) or measure the cosmological properties of our universe by (see, e.g. Schutz, 1986; Abac et al., 2025g). These methods typically perform inference on a hierarchical population model using the Gravitational-Wave Transient Catalog (GWTC) posteriors as input data (Thrane & Talbot, 2019). On the other hand, considering individual events, studies can be conducted using the posterior samples to examine the inferences under alternative assumptions about the waveform model, prior distribution, or data. Underlying all these methods is the technique known as resampling, which utilises a set of weights, constructed from the ratio of the new to the old posterior density, to generate a new set of samples under some changed assumptions. Resampling is an alternative to a re-analysis (i.e. applying stochastic sampling to estimate the posterior afresh) which can be computationally prohibitive.
@@
Two types of resampling that are in common use: RS and IS. For RS (discussed further in Section˜3.1), samples are accepted into the new set in proportion to their weight. Meanwhile, IS (discussed further in Section˜3.2) is a broad term, often without a consensus definition, but generally refers to the process of using the weights to determine the importance of the samples. In some use cases, this involves using the weights and samples together (e.g. to create a weighted histogram). Alternatively, some use cases apply multinomial sampling to produce a new sample set according to the probabilities provided by the weights (either with or without replacement). In this work, we refer to the former as IS and the latter as multinomial-IS.
@@
Resampling techniques have long been used in the field. For example, both Markov chain Monte Carlo (MCMC) and nested sampling utilise resampling to produce sample sets; Rover et al. (2006) used multinomial-IS to initialise a MCMC sampler; Payne et al. (2019) demonstrated the use of IS to utilise higher-order gravitational-wave models; Payne et al. (2020) extended this to marginalize over the calibration model (and this was implemented in Abbott et al. (2024)); Baka et al. (2025) explored RS to correct for calibration misspecification; Tiwari (2018) and Talbot et al. (2019) used IS to estimate the binary black hole (BBH) population sensitivity; RS is used to apply a cosmological prior to LVK analyses (Abac et al., 2025d), and an astrophysically-motivated prior (Chattopadhyay et al., 2024; Abac et al., 2025e); Dax et al. (2023) uses IS with a proposal distribution provided by neural posterior estimation (but generally the technique can be applied to any pre-computed sample set with a tractable posterior); Williams et al. (2025) investigated sequential-Monte Carlo approaches, an extension of IS; and finally, Hourihane & Chatziioannou (2025), use IS changing the data in the likelihood to study the impact of transient non-Gaussian noise. There are many more examples besides this in the literature. However, we caution that the terms are often confused and there are additional terms such as “reweighting” that we choose to avoid since this is used to describe both RS and IS.
@@
The goal of this work is to provide a simple guide to the application of resampling, along with a discussion of reconstructing the likelihood and prior distribution. We begin in Section˜2 by discussing how the likelihood and prior can be calculated from the released posterior samples, including a comparison with the original calculations stored in data files. We also provide estimates of the accuracy achievable given practical hardware and software considerations. Then, in Section˜3, we go on to introduce and compare the methodology of resampling, including RS and IS and introduce Pareto-smoothing as a means to improve the efficiency. In Section˜4, we provide examples using publically available data sets and applying resampling to illustrate the possibilities. Finally, we conclude with a discussion in Section˜5. Readers may find it useful to refer to the data release associated with this work (Ashton, 2025), which contains worked examples.
@@ 2 Reconstructing the likelihood and prior
In this Section, we will detail how to reconstruct the likelihood and prior for samples drawn from the posterior distribution presented in the digital GWTC. Specifically, we consider existing analyses of gravitational-wave signals using the Bilby inference library and then packaged using PESummary. Examples of such files can be found as part of the data released with GWTC-4.0 (LIGO Scientific Collaboration et al., 2025a).
Each file stores a set of posterior samples $\{\theta_{i}\}$. Additionally, users will find columns corresponding to the log-likelihood and log-prior (in all cases, these are natural-logarithm). We refer to these as the stored values and denote them with a sub-script $S$ (e.g. $\ln\mathcal{L}_{S}$ and $\ln\pi_{S}$) to differentiate them from the reconstructed values that we denote with a sub-script $R$ (e.g. $\ln\mathcal{L}_{R}$ and $\ln\pi_{R}$). For accurate reconstruction of the log-likelihood and log-prior for a sample, it is important to match the original computation. Any differences can manifest as either systematic shifts (which could entirely bias any subsequent resampling) or random shifts (which, if sufficiently small, may be negligible). In the following, we address each point in turn.
@@ 2.1 Computing environment
First, the versions of the software should match those used in the analysis: this is because changes to the software may reflect changes to the underlying assumptions. For example, if the default assumptions about data processing change between versions of the software, this will induce systematic shifts in the likelihood. Similarly, if the waveform model is not identical (including specific options hardcoded into the environment), this can also induce systematic and random shifts in the likelihood. Information about the software packages used in the analysis is packaged as part of the PESummary data release (see notebooks).
@@
Second, for exact reconstruction, the hardware should also be identical. This is because the underlying hardware determines the implementation of floating-point arithmetic: differences in this implementation will change the byte-level results. There may also be platform-specific optimisation of linear algebra libraries that can differ even if the hardware is identical. Comparing an analysis run on an Intel Xeon Processor (Skylake) and then reconstructed on an Apple M3 Pro, we find the distributions of log-likelihoods are centred on zero but with a standard deviation of $\approx 3\times 10^{-6}$. Meanwhile, when they are reconstructed on identical hardware, we find the log-likelihoods can be reconstructed up to floating point precision. However, typically, it is not possible to match the hardware used in the analysis. But, as we will show later, differences of less than 1 part in a million are sufficiently small that they have a negligible impact on the capacity to resample the posterior.
@@ 2.2 Likelihood
To reconstruct the likelihood, Eq.˜3, for a given sample, we must match the data, PSD, waveform model, and likelihood configuration. In the attached notebook, we demonstrate how this can be achieved, taking the settings from the configuration file packaged as part of the data release. In the following, we provide a discussion of specific elements.
@@
Data – The raw strain data used in LVK analyses is available from the Gravitational Wave Open Science Center (GWOSC) at https://gwosc.org/data/ and available either at the original sampling frequency of $16382$ Hz or downsampled to $4192$ Hz. While there is not expected to be any signal content above a few thousand Hertz, for accurate reconstruction, it is important to start from the data sampled at $16382$ Hz since the details of the downsampling algorithm do impact the entire data set: using the pre-downsampled data can result in significant deviations between the reconstructed and original likelihoods. For some events, the raw data is affected by transient non-Gaussian noise artefacts known as glitches (Nuttall, 2018; Glanzer et al., 2023; Soni et al., 2025). In many cases, the glitches can be modelled and removed from the data using either a modelled Bayesian approach (BayesWave: Pankow et al., 2018; Cornish et al., 2021; Chatziioannou et al., 2021; Hourihane et al., 2022), or a linear noise subtraction based on auxiliary witness channels (Davis et al., 2022). Therefore, to recreate the likelihood, the glitch-subtracted data should be used (see, e.g. LIGO Scientific Collaboration et al., 2025b, for glitch-subtracted data for events from the first part of the fourth observing run).
@@
Downsampling – For LVK analyses, Bilby uses the ResampleTimeSeries downsampling routines implemented in LALSuite which implements a time-domain Butterworth filter (Butterworth et al., 1930). This filter was selected to match the downsampling performed by the tools used to generate the PSD (discussed shortly). The specific choice of sampling frequency is stored in the configuration file for the Bilby analysis.
@@
Windowing – The strain data used for inference is a subset of a longer time series of strain taking during the observation run. Typically, a short section of data is cut out, containing the observable signal and sufficient padding on either side. As a result, this time series is not periodic, which is required to approximately diagonalise the noise covariance matrix (see the discussion in Talbot et al., 2025). Therefore, a window factor is applied, gradually tapering the time series to zero at either end. Following standard methods (Abac et al., 2025d), a Tukey window (Harris, 2005) is typically applied with a roll-off determined by a factor determining the fraction of the window that is tapered. Before passing the data to likelihood, this window should be applied to the time-domain strain.
@@
PSD – The GWTC analyses use BayesWave (Cornish & Littenberg, 2015; Littenberg & Cornish, 2015; Cornish et al., 2021; Gupta & Cornish, 2024) to estimate the on-source PSD for each detector. This is packaged within the data release and can be extracted and provided to the Bilby likelihood. Note that the data pre-processing should be identical for the analysis data and PSD construction.
@@
Explicit marginalization – The likelihood used for inference, at base, is the Whittle likelihood as given in Eq.˜3. However, Bilby implements several forms of explicit marginalisation, and these are routinely used to improve the convergence properties of the samplers (Abac et al., 2025d). As further described in Thrane & Talbot (2019); Romero-Shaw et al. (2020), after sampling Bilby then reconstructs the full posterior distribution in post-processing. Therefore, it is important to differentiate between the marginalized likelihood which is used for sampling and the non-marginalised likelihood, which corresponds to the full posterior distribution. In the attached notebooks, we show how both of these can be calculated. In addition, we show that the log-likelihood that is stored in the data release corresponds to the marginalized likelihood. The not-marginalised likelihood is not stored within the packaged data release, but can be calculated from the matched-filter and optimal signal-to-noise ratios that are stored (as shown in the data release).

### v2

@@ 1 Introduction
Gravitational wave observations of merging black holes and neutron stars are now routinely made by interferometric gravitational wave detectors. There are currently four operational kilometre-scale interferometes: Laser Interferometer Gravitational-Wave Observatory (LIGO: Aasi and others, 2015), Virgo (Acernese and others, 2015), and KAGRA (Akutsu and others, 2021). Their raw output is a time series of calibrated strain, representing the relative difference in arm lengths for each instrument. The –Virgo–KAGRA (LVK) Collaborations publish this data (Abac and others, 2025f) in addition to the Gravitational-Wave Transient Catalog (GWTC: Abac and others, 2025a), containing a list of candidate signals and measurements of their Bayesian posterior distributions (Abac and others, 2025e).
The posterior distribution (see Thrane and Talbot, 2019, for a discussion on why a Bayesian framework is used) is defined as the probability density of a set of source parameters $\theta$ conditional on the data $d$ and model $M$ (which implicitly defines the parameters) and can be calculated up to a normalisation factor from
@@
$\mathcal{Z}(d|M)=\int\mathcal{L}(d|\theta,M)\pi(\theta|M)\,d\theta\,.$ (2)
@@
The LVK analysis methodology for approximating Eq.˜1 for observed signals is described in detail in Section 5 of Abac and others (2025c). In brief, for transient gravitational-wave data analysis, analysis is performed on a time series of consecutive strain data observations, which we denote by $d$ and define its duration to be $T$ and sampling frequency $f_{s}$. Inference is performed on the frequency series $\tilde{d}$ generated by applying a fast Fourier transform and dividing by $f_{s}$ (a normalisation choice; see Thrane and Talbot (2019)). For signals lasting up to a few hundred seconds, we assume the non-astrophysical noise to be generated by a stationary additive coloured Gaussian noise process with a power spectral density (PSD) $P$ such that the likelihood for the $j$th frequency bin is given by
$\mathcal{L}(d_{j}|\theta,M)=\frac{1}{2\pi P_{j}}\exp\left(-\frac{2}{T}\frac{|\tilde{d}_{j}-\tilde{\mu}_{j}(\theta)|^{2}}{P_{j}}\right)\,$, (3)
@@
where $\tilde{\mu}_{j}(\theta)$ is the frequency-domain astrophysical signal model projected onto the response of the detector. The full likelihood is then calculated from the product over the $j$ frequency bins, i.e.
$\mathcal{L}(d|\theta,M)=\prod_{j}\mathcal{L}(d_{j}|\theta,M)\,.$ (4)
In practise, this is computed as the sum in log-space for numerical stability.
@@
Typically, $\tilde{\mu}_{j}(\theta)$ is a waveform approximant that models the inspiral–merger–ringdown of the compact binary coalescence (CBC) signal utilising either phenomenological fits (Ajith and others, 2007), the effective one-body approach (Buonanno and Damour, 1999, 2000), or numerical-relativity surrogates (Blackman et al., 2017). To date, Bilby uses the LALSuite (LIGO Scientific Collaboration et al., 2018; Wette, 2020) library to access these waveforms.
@@
Meanwhile, the prior distribution $\pi(\theta|M)$ is either generally weakly astrophysically motivated or informed by the observed population of CBC sources (Abac and others, 2025c).
@@
Eq.˜1 is not solvable in closed form for the CBC likelihood and prior, and the distribution is typically non-Gaussian, multimodal, and contains strong curved degeneracies. Therefore, computational Bayesian inference methods such as Markov chain Monte Carlo (MCMC; Hastings, 1970) and nested sampling (Skilling, 2006) are used with post-processing to produce a set of equally-weighted posterior samples $\theta_{i}\sim p(\theta|d,M)$ and an estimate of the evidence $\mathcal{Z}(d|M)$. To date, the LVK has used LALInference (Veitch and others, 2015), RIFT (Pankow et al., 2015; Lange and others, 2017; Wysocki et al., 2019), and Bilby (Ashton and others, 2019; Romero-Shaw and others, 2020), with an adaptation of the Dynesty (Speagle, 2020) nested sampling implementation, to produce posterior samples. In this work, we will consider the LVK posteriors produced by Bilby and packaged using PESummary (Hoy and Raymond, 2021). This is motivated by the ease of use in reconstructing the posteriors using Bilby (part of its design philosophy) and the wide applicability: Bilby posteriors have been estimated for all signals above threshold for parameter estimation from the first observing run until the most recent release, see GWTC-2.1 (Abbott and others, 2024), GWTC-3.0 (Abbott and others, 2023), and GWTC-4.0 (Abac and others, 2025e).
@@
The posterior samples from a given observed signal form a data set that can be used for further analyses. For example, taking samples from a set of highly probable signals, the samples form an input data set that can be used to constrain the astrophysical distribution of merging black holes (see, e.g. Talbot and Thrane, 2018; Abac and others, 2025d) or measure the cosmological properties of our universe (see, e.g. Schutz, 1986; Abac and others, 2025b). These methods typically perform inference on a hierarchical population model using the Gravitational-Wave Transient Catalog (GWTC) posteriors as input data (Thrane and Talbot, 2019). On the other hand, considering individual events, studies can be conducted using the posterior samples to examine the inferences under alternative assumptions about the waveform model, prior distribution, or data. Underlying all these methods is the technique known as resampling, which utilises a set of weights, constructed from the ratio of the new to the old posterior density, to generate a new set of samples under some changed assumptions. Resampling is an alternative to a re-analysis (i.e. applying stochastic sampling to estimate the posterior afresh), which can be computationally prohibitive.
@@
Two types of resampling are in common use: RS and IS. For RS (discussed further in Section˜3.1), samples are accepted into the new set in proportion to their weight. Meanwhile, IS (discussed further in Section˜3.2) is a broad term, often without a consensus definition, but generally refers to the process of using the weights to determine the importance of the samples. In some use cases, this involves using the weights and samples together (e.g. to create a weighted histogram). Alternatively, some use cases apply multinomial sampling to produce a new sample set according to the probabilities provided by the weights (either with or without replacement). In this work, we refer to the former as IS and the latter as multinomial-IS.
@@
Resampling techniques have long been used in the field. For example, both Markov chain Monte Carlo (MCMC) and nested sampling utilise resampling to produce sample sets; Rover et al. (2006) used multinomial-IS to initialise a MCMC sampler; Payne et al. (2019) demonstrated the use of IS to utilise higher-order gravitational-wave models; Payne et al. (2020) extended this to marginalise over the calibration model (and this was implemented in Abbott and others (2024)); Baka et al. (2025) explored RS to correct for calibration misspecification; Tiwari (2018) and Talbot et al. (2019) used IS to estimate the binary black hole (BBH) population sensitivity; RS is used to apply a cosmological prior to LVK analyses (Abac and others, 2025c), and an astrophysically-motivated prior (Chattopadhyay et al., 2024; Abac and others, 2025e); Dax et al. (2023) uses IS with a proposal distribution provided by neural posterior estimation (but generally the technique can be applied to any pre-computed sample set with a tractable posterior); Williams et al. (2025) investigated sequential-Monte Carlo approaches, an extension of IS; and finally, Hourihane and Chatziioannou (2025), use IS changing the data in the likelihood to study the impact of transient non-Gaussian noise. There are many more examples besides this in the literature. However, we caution that the terms are often confused and there are additional terms such as “reweighting” that we choose to avoid since this is used to describe both RS and IS.
@@
The goal of this work is to provide a simple guide to the application of resampling, along with a discussion of reconstructing the likelihood and prior distribution. We begin in Section˜2 by discussing how the likelihood and prior can be calculated from the released posterior samples, including a comparison with the original calculations stored in data files. We also provide estimates of the accuracy achievable given practical hardware and software considerations. Then, in Section˜3, we go on to introduce and compare the methodology of resampling, including RS and IS and introduce Pareto-smoothing as a means to improve the efficiency. In Section˜4, we provide examples using publicly available data sets and applying resampling to illustrate the possibilities. Finally, we conclude with a discussion in Section˜5. Readers may find it useful to refer to the data release associated with this work (Ashton, 2026), which contains worked examples.
@@ 2 Reconstructing the likelihood and prior
In this Section, we will detail how to reconstruct the likelihood and prior for samples drawn from the posterior distribution presented in the digital GWTC. Specifically, we consider existing analyses of gravitational-wave signals using the Bilby inference library and then packaged using PESummary. Examples of such files can be found as part of the data released with GWTC-4.0 (LIGO Scientific Collaboration et al., 2025b).
Each file stores a set of posterior samples $\{\theta_{i}\}$. Additionally, users will find columns corresponding to the log-likelihood and log-prior (in all cases, these are natural logarithms). We refer to these as the stored values and denote them with a sub-script $S$ (e.g. $\ln\mathcal{L}_{S}$ and $\ln\pi_{S}$) to differentiate them from the reconstructed values that we denote with a sub-script $R$ (e.g. $\ln\mathcal{L}_{R}$ and $\ln\pi_{R}$). For accurate reconstruction of the log-likelihood and log-prior for a sample, it is important to match the original computation. Any differences can manifest as either systematic shifts (which could entirely bias any subsequent resampling) or random shifts (which, if sufficiently small, may be negligible). In the following, we address each point in turn.
@@ 2.1 Computing environment
First, the versions of the software should match those used in the analysis: this is because changes to the software may reflect changes to the underlying assumptions. For example, if the default assumptions about data processing change between versions of the software, this will induce systematic shifts in the likelihood. Similarly, if the waveform model is not identical (including specific options hardcoded into the environment), this can also induce systematic and random shifts in the likelihood. Information about the software packages used in the analysis is packaged as part of the PESummary results files (see the data release for details on how to access the information).
@@
Second, for exact reconstruction, the hardware should also be identical. This is because the underlying hardware determines the implementation of floating-point arithmetic: differences in this implementation will change the byte-level results. There may also be platform-specific optimisation of linear algebra libraries that can differ even if the hardware is identical. Comparing an analysis run on an Intel Xeon Processor (Skylake) and then reconstructed on an Apple M3 Pro, we find the distributions of log-likelihoods are centred on zero but with a standard deviation of $\approx 3\times 10^{-6}$. Meanwhile, when they are reconstructed on identical hardware, we find the log-likelihoods can be reconstructed up to floating-point precision. However, typically, it is not possible to match the hardware used in the analysis. But, as we will show later, differences of less than 1 part in a million are sufficiently small that they have a negligible impact on the capacity to resample the posterior.
@@ 2.2 Likelihood
To reconstruct the likelihood, Eq.˜3, for a given sample, we must match the data, PSD, waveform model, and likelihood configuration. In the data release, we demonstrate how this can be achieved, taking the settings from the configuration file packaged as part of the data release. In the following, we provide a discussion of specific elements.
The likelihood ratioEq.˜3 is the likelihood of the data under the model comprised of an additive signal and noise. However, we can see that if $\mu_{j}=0\;\forall\;j$, i.e. there is no signal, just noise, then we can also introduce $\mathcal{L}(d|N)$, the noise likelihood. Since the noise likelihood is $\theta$-independent, it is common during sampling to use the likelihood ratio
$\Lambda(d|\theta,M)=\frac{\mathcal{L}(d|\theta,M)}{\mathcal{L}(d|N)}\,$, (5)
rather than the likelihood itself. While this choice does not change the parameter estimates, it is often used because the evidence estimate obtained (e.g. from nested sampling) is the signal vs noise Bayes factor, i.e. ${\mathcal{B}}=\mathcal{Z}(d|M)/\mathcal{Z}(d|N)$, which is easier to interpret as a diagnostic during sampling rather than $\mathcal{Z}{(d|M)}$ itself.
When reconstructing the likelihood, it is important to distinguish between the likelihood and the likelihood ratio. For resampling, one only needs to be consistent However, evidence estimates computed from the likelihood will yield measurements of the signal evidence. In contrast, those computed from the likelihood ratio will yield measurements of the signal vs noise Bayes factor. Bilby can compute either as demonstrated in the data release.
@@
Data – The raw strain data used in LVK analyses is available from the Gravitational Wave Open Science Center (GWOSC) at https://gwosc.org/data/ and available either at the original sampling frequency of $16384$ Hz or downsampled to $4096$ Hz. While there is not expected to be any signal content above a few thousand Hertz, for accurate reconstruction, it is important to start from the data sampled at $16382$ Hz since the details of the downsampling algorithm do impact the entire data set: using the pre-downsampled data can result in significant deviations between the reconstructed and original likelihoods. For some events, the raw data is affected by transient non-Gaussian noise artefacts known as glitches (Nuttall, 2018; Glanzer and others, 2023; Soni and others, 2025). In many cases, the glitches can be modelled and removed from the data using either a modelled Bayesian approach (BayesWave: Pankow and others, 2018; Cornish et al., 2021; Chatziioannou et al., 2021; Hourihane et al., 2022), or a linear noise subtraction based on auxiliary witness channels (Davis et al., 2022). Therefore, to recreate the likelihood, the glitch-subtracted data should be used (see, e.g. LIGO Scientific Collaboration et al., 2025a, for glitch-subtracted data for events from the first part of the fourth observing run).
@@
Downsampling – For LVK analyses, Bilby uses the ResampleTimeSeries downsampling routines implemented in LALSuite which implements a time-domain Butterworth filter (Butterworth and others, 1930). This filter was selected to match the downsampling performed by the tools used to generate the PSD (discussed shortly). The specific choice of sampling frequency is stored in the configuration file for the Bilby analysis.
@@
Windowing – The strain data used for inference is a subset of a longer time series of strain taken during the observation run. Typically, a short section of data is cut out, containing the observable signal and sufficient padding on either side. As a result, this time series is not periodic, which is required to approximately diagonalise the noise covariance matrix (see the discussion in Talbot et al., 2025). Therefore, a window factor is applied, gradually tapering the time series to zero at either end. Following standard methods (Abac and others, 2025c), a Tukey window (Harris, 2005) is typically applied with a roll-off determined by a factor determining the fraction of the window that is tapered. Before passing the data to likelihood, this window should be applied to the time-domain strain.
@@
PSD – The GWTC analyses use BayesWave (Cornish and Littenberg, 2015; Littenberg and Cornish, 2015; Cornish et al., 2021; Gupta and Cornish, 2024) to estimate the on-source PSD for each detector. This is packaged within the data release and can be extracted and provided to the Bilby likelihood. Note that the data pre-processing should be identical for the analysis data and PSD construction.
@@
Explicit marginalization – The likelihood used for inference, at base, is the Whittle likelihood as given in Eq.˜3. However, Bilby implements several forms of explicit marginalisation, and these are routinely used to improve the convergence properties of the samplers (Abac and others, 2025c). As further described in Thrane and Talbot (2019); Romero-Shaw and others (2020), after sampling Bilby then reconstructs the full posterior distribution in post-processing. Therefore, it is important to differentiate between the marginalised likelihood, which is used for sampling and the non-marginalised likelihood, which corresponds to the full posterior distribution. In the data release, we show how both of these can be calculated. In addition, we show that the log-likelihood that is stored in the data release corresponds to the marginalised likelihood. The not-marginalised likelihood is not stored within the packaged data release, but can be calculated from the matched-filter and optimal signal-to-noise ratios that are stored (as shown in the data release).